OpenStudy (anonymous):

Please help! I've been stuck on these questions for weeks now! Share 5 cookies with 4 people. How much cookie should each person get? Express your answer as a series. Share 2 cookies with 3 people. How much cookie should each person get? Express your answer as a series.

OpenStudy (thomas5267):

What series do they want? I thought series is the sum of a sequence.

OpenStudy (anonymous):

The first is a finite geometric series. The second is an infinite series. I thought the same thing. That's why I've been so stuck on this

OpenStudy (thomas5267):

So I have to express \(\dfrac{5}{4}\) as a infinite geometric series?

OpenStudy (anonymous):

Finite

OpenStudy (thomas5267):

\[ \begin{align*} r+r^2&=\frac{5}{4}\\ r^2+r-\frac{5}{4}&=0\\ r&=\frac{1}{2}(\sqrt{6}-1) \end{align*} \] The geometric sequence is \(\left(\frac{1}{2}(\sqrt{6}-1)\right)^n\) and the sum of first two terms is \(\dfrac{5}{4}\).

OpenStudy (thomas5267):

Just create a geometric series that works. You could have at least one geometric series with n-terms which sums to \(\dfrac{5}{4}\) for every n I think.

OpenStudy (anonymous):

I'm sorry, I'm still confused..

OpenStudy (anonymous):

I think they explained enough.

OpenStudy (phi):

The sum of an infinite geometric series is (see https://en.wikipedia.org/wiki/Geometric_series#Formula) is \[S=a+ar+ar^2+ar^3+...=\sum_{k=0}^{\infty}ar^k= \frac{ a }{ 1-r }\] we want this to add up to 5/4 Let's set a=1 (we can choose any number except 0, so keep it simple) we now want the sum to be equal to 5/4 so it must be that \[ \frac{ 1 }{ 1-r }= \frac{5}{4}\] now solve for r. I would cross-multiply \[ 4 = 5(1-r) \\ 4= 5 -5r\] add -4 to both sides \[ 0 = 5 - 4 - 5r \\ 0 = 1 - 5r\] add +5r to both sides \[ 5r= 1 \] divide both sides by 5 \[ r= \frac{1}{5} \] that means the infinite series with a=1 and r= 1/5 is: \[ 1 + \frac{1}{5} + \frac{1}{25} + \frac{1}{125} + ... \]